# FastTracker: Real-Time and Accurate Visual Tracking

### v1

@@ Abstract
Conventional multi-object tracking (MOT) systems are predominantly designed for pedestrian tracking and often exhibit limited generalization to other object categories. This paper presents a generalized tracking framework capable of handling multiple object types, with a particular emphasis on vehicle tracking in complex traffic scenes. The proposed method incorporates two key components: (1) an occlusion-aware re-identification mechanism that enhances identity preservation for heavily occluded objects, and (2) a road-structure-aware tracklet refinement strategy that utilizes semantic scene priors—such as lane directions, crosswalks, and road boundaries—to improve trajectory continuity and accuracy. In addition, we introduce a new benchmark dataset comprising diverse vehicle classes with frame-level tracking annotations, specifically curated to support evaluation of vehicle-focused tracking methods. Extensive experimental results demonstrate that the proposed approach achieves robust performance on both the newly introduced dataset and several public benchmarks, highlighting its effectiveness in general-purpose object tracking. While our framework is designed for generalized multi-class tracking, it also achieves strong performance on conventional benchmarks, with HOTA scores of 66.4 on MOT17 and 65.7 on MOT20 test sets. Code and Benchmark are available: github.com/Hamidreza-Hashempoor/FastTracker, huggingface.co/datasets/Hamidreza-Hashemp/FastTracker-Benchmark.
@@
Visual Tracking
@@ I Introduction
Multi-object tracking (MOT) plays a critical role in computer vision applications such as intelligent surveillance and autonomous driving. Despite significant progress, it remains a challenging problem due to factors like target similarity, frequent occlusions, and the continuous entry and exit of objects from the scene [1]. A widely adopted approach to address MOT is the tracking-by-detection paradigm [2], where object detectors first identify candidate targets in each frame, and a separate tracking module associates these detections with existing trajectories. This association is typically formulated as a matching problem, relying on computed similarity measures between current detections and existing tracklets. Algorithms such as the Hungarian method [3], have been employed to solve this assignment task effectively. Detection algorithms also provide a confidence score for each bounding box, which reflects the likelihood of a detection being valid. In general, high-confidence detections are expected to correspond to true positives, while lower scores often indicate false positives.
Figure 1: Comparison of multi-object tracking performance using five key metrics: HOTA, MOTA, IDF1, DetA, and AssA on MOT17 test set. Each axis is normalized to its respective scale range to enable a direct visual comparison, highlighting trade-offs between detection accuracy and association quality.
@@ II Related Work
III Approach
To improve data association in multi-object tracking, we adopt a straightforward yet powerful strategy that avoids discarding low-confidence detections outright. Rather than filtering out detections based solely on confidence thresholds—as done in many earlier approaches—we divide the detections into high- and low-confidence groups similar to Byte tracker. This separation enables more flexible and robust tracking: high-confidence detections provide strong initial associations, while selectively using low-confidence detections helps recover missed targets and reduce identity switches. This two-stage design improves overall recall without increasing false positives, especially in crowded or ambiguous scenes.
However, in practice, severe occlusions and dense traffic still cause association failures. To mitigate this, we introduce explicit occlusion detection metrics based on spatial overlap, allowing occluded targets to be marked and handled even without detections. We also incorporate environment-aware cues—such as road directionality, street layout, and pedestrian zones—to constrain and refine motion trajectories, preventing unrealistic movements during occlusion. Additionally, we revise the tracklet initialization and removal logic by applying more selective confirmation and deletion policies. These changes reduce spurious identities, improve recovery of true targets, and ensure occluded objects are handled separately. Together, these enhancements significantly improve tracking robustness in complex, multi-class urban scenes. The complete FastTrack algorithm is summarized below.
Figure 4: Illustration of DampenVelocity and EnlargeBox. Top row shows standard tracking where fast motion and tight boxes lead to ID switches. Bottom row shows how velocity damping and box enlargement help maintain consistent identities during occlusion.
@@ III Approach
Motion Prediction.   For each tracklet $t\in\mathcal{T}$, we estimate its future state $\hat{x}_{t}$ using a class-aware Kalman filter. The motion model parameters are selected based on the object class: vehicles such as cars or motorcycles are allowed higher velocities and acceleration bounds, whereas pedestrians are modeled with smoother and slower dynamics. This enables more realistic trajectory predictions, especially in cases of temporary occlusion or detection dropout.
@@
Direction and ROI Constraint.   To prevent tracklets from drifting into implausible directions, we retrieve a region $R$ around the predicted state $\hat{x}_{t}$ using a region lookup over the environment map $\mathcal{M}$. Each region encodes semantic layout information—such as drivable roads or pedestrian paths—and is associated with a motion cone reflecting the allowed directionality. The cone angle $\theta$ is computed geometrically based on the polygonal structure of the region: we identify the entrance and exit edges (defined by prior knowledge of scene flow), then construct two diagonals connecting each corner of the entrance edge to its diagonally opposite point on the exit edge. The angle $\theta$ is defined as the angle between these two crossing diagonals, representing the allowable motion range. The function $\mathrm{ProjectToCone}(\hat{x}_{t},\mathrm{cone}(R))$ projects the predicted position into this direction-constrained cone. We illustrate these concepts in the accompanying figure 3 using a one-sided road, though the same logic extends to other region types such as two-sided roads or crosswalks.
@@
Association.   We adopt a two-stage association strategy similar to Byte Tracker. In the first stage, high-confidence detections $\mathcal{D}_{\mathrm{high}}$ are matched with active tracklets $\mathcal{T}$ using IoU-based association, yielding matches $(\mathcal{T}_{1},\mathcal{D}_{1})$. Unmatched tracklets and low score detections, $\mathcal{T}_{\mathrm{remain}}$ and $\mathcal{D}_{\mathrm{low}}$, are then passed to the second stage, where $\mathcal{T}_{\mathrm{remain}}$ is matched with low-confidence detections $\mathcal{D}_{\mathrm{low}}$ to recover hard cases $\mathcal{T}_{2}$ and we seclude them from $\mathcal{T}_{\mathrm{remain}}$. Finally, the updated active set becomes $\mathcal{T}=\mathcal{T}_{1}\cup\mathcal{T}_{2}$.
@@
Occlusion Handling.   Unmatched tracklets $\mathcal{T}_{\mathrm{remain}}$ are checked for occlusion by measuring spatial overlap with active tracklets $\mathcal{T}$. If the center-proximity score $\mathrm{CP}(t,t^{\prime})$, calculated by IoU, exceeds a threshold $\mathrm{CP}_{\min}$, the tracklet $t$ is considered occluded and added to $\mathcal{T}_{\mathrm{occ}}$. For each occluded tracklet, we mark it as occluded, dampen its velocity to avoid propagation in the direction of the occluding object (which often causes drift), and enlarge its bounding box by 10–20% (depending on object class) to improve future re-identification as detailed in Figure 4. Occluded tracklets are temporarily removed from further association steps until they reappear. The effectiveness of bounding box enlargement and velocity damping is illustrated in Figure 5
@@ V Experimental Setup
Implementation Details.    For detection, we adopt YOLOX due to its effective balance between speed and accuracy. The detector’s classification and localization heads are trained following the best practices established in prior work [5, 35]. At inference time, we apply a non-maximum suppression (NMS) threshold of 0.75. Tracklet association uses an IoU threshold of 0.5, and exponential moving average (EMA) smoothing is applied with a momentum coefficient $\alpha$ of 0.8. We set $\mathrm{CP}_{\min}$ to 0.7, the initialization overlap threshold $K_{init}$ to 0.8, and the occlusion tolerance window $T_{occ}$ to 30 frames. Detection confidence thresholds for classification are $\tau_{low}=0.2$ and $\tau_{high}=0.65$. The region of interest $\mathcal{M}$ is user-configurable, and if direction information is provided, the tracker adapts accordingly.
@@ VI-A Ablation Studies:
Post Processing: To evaluate the contribution of the post-processing stage, we analyze the impact of incorporating the GSP, which outperforms conventional interpolation techniques [37], and global linking (G-Link) on the final tracking performance across three datasets. Importantly, all results reported here are based on the full online system, which includes our proposed deletion/initialization policies and occlusion handling mechanisms. The post-processing modules are applied offline after generating the initial tracks. As shown in Table V, enabling either GSP or G-Link individually brings consistent improvements over the base output, with G-Link slightly outperforming GSP in terms of HOTA across all datasets.
@@ VI-B Benchmark Evaluation
FastTracker Benchmark:    Table X presents the results on the FastTracker Benchmark, a challenging internal benchmark designed to assess robustness in dense tracking scenarios. FastTracker achieves the best MOTA (63.8), IDF1 (79.2) and HOTA (61.0), surpassing FeatureSORT and ByteTrack in association accuracy. It also achieves the lowest number of identity switches (251), demonstrating strong consistency across frames. While FastTracker leads in MOTA (63.8), it also shows a better balance between detection and identity preservation, outperforming StrongSORT and FeatureSORT by +1.3 and +3 HOTA, respectively, and significantly improving IDF1 by up to +1.8. These results highlight FastTracker’s strength in maintaining accurate identity tracking under crowded conditions.
@@
VII Discussion
While FastTracker demonstrates strong performance across public and internal benchmarks, some limitations remain. Currently, the system relies on manually defined ROI regions and cone direction constraints, which must be specified using exactly four edges. This rigid configuration may limit deployment in complex or dynamic scenes where such annotations are impractical or insufficient. As a promising direction for future work, recent advances in semantic segmentation and scene understanding can be leveraged to enable automatic detection of road boundaries, crosswalks, and other contextual cues, removing the need for manual setup. Moreover, extending the system to support arbitrary polygonal ROIs or more flexible directional constraints would allow broader applicability in real-world environments such as intersections, roundabouts, and multi-lane roads.
VIII Conclusion

### v3

@@ Abstract
Conventional multi-object tracking (MOT) systems are predominantly designed for pedestrian tracking and often exhibit limited generalization to other object categories. This paper presents a generalized tracking framework capable of handling multiple object types, with a particular emphasis on vehicle tracking in complex traffic scenes. The proposed method incorporates two key components: (1) an occlusion-aware re-identification mechanism that enhances identity preservation for heavily occluded objects, and (2) a road-structure-aware tracklet refinement strategy that utilizes semantic scene priors—such as lane directions, crosswalks, and road boundaries—to improve trajectory continuity and accuracy. In addition, we introduce a new benchmark dataset comprising diverse vehicle classes with frame-level tracking annotations, specifically curated to support evaluation of vehicle-focused tracking methods. Extensive experimental results demonstrate that the proposed approach achieves robust performance on both the newly introduced dataset and several public benchmarks, highlighting its effectiveness in general-purpose object tracking. While our framework is designed for generalized multi-class tracking, it also achieves strong performance on conventional benchmarks, with HOTA scores of 66.4 on MOT17 and 65.7 on MOT20 test sets. github.com/Hamidreza-Hashempoor/FastTracker, huggingface.co/datasets/Hamidreza-Hashemp/FastTracker-Benchmark.
@@
Multi Object Tracking, Visual Tracking
@@ I Introduction
Multi object tracking (MOT) plays a critical role in computer vision applications such as intelligent surveillance and autonomous driving. Despite significant progress, it remains a challenging problem due to factors like target similarity, frequent occlusions, and the continuous entry and exit of objects from the scene [1]. A widely adopted approach to address MOT is the tracking-by-detection paradigm [2], where object detectors first identify candidate targets in each frame, and a separate tracking module associates these detections with existing trajectories. This association is typically formulated as a matching problem, relying on computed similarity measures between current detections and existing tracklets. Algorithms such as the Hungarian method [3], have been employed to solve this assignment task effectively. Detection algorithms also provide a confidence score for each bounding box, which reflects the likelihood of a detection being valid. In general, high-confidence detections are expected to correspond to true positives, while lower scores often indicate false positives.
Figure 1: Comparison of multi-object tracking performance using five key metrics: HOTA, MOTA, IDF1, DetA, and AssA on MOT17 test set. Each axis is within the specified range, highlighting trade-offs between detection accuracy and association quality.
@@ II Related Work
Figure 3: Overall pipeline of Online FastTracker. Kalman prediction is refined with region constraints (ProjectToCone, ClampToROI), followed by two-stage association using high- and low-confidence detections. Occlusion handling (DampenVelocity, EnlargeBox), deletion, and initialization ensure stable and consistent tracklet management.
@@
III Background
In tracking-by-detection frameworks such as ByteTrack [5], data association is performed in two stages to balance recall and precision. Let $\mathcal{T}$ denote the set of active tracklets at the current frame, and let the detector provide two groups of detections: $\mathcal{D}_{\mathrm{high}}$ for high-confidence detections (above threshold $\tau_{\mathrm{high}}$) and $\mathcal{D}_{\mathrm{low}}$ for low-confidence detections (between $\tau_{\mathrm{low}}$ and $\tau_{\mathrm{high}}$). In the first stage, $\mathcal{D}_{\mathrm{high}}$ is matched with $\mathcal{T}$ using IoU-based similarity, yielding matched pairs $(\mathcal{T}_{1},\mathcal{D}_{1})$. Unmatched tracklets and detections are carried forward as $\mathcal{T}_{\mathrm{remain}}$ and $\mathcal{D}_{\mathrm{remain}}$, respectively. In the second stage, the unmatched tracklets $\mathcal{T}_{\mathrm{remain}}$ are further matched with $\mathcal{D}_{\mathrm{low}}$, producing additional associations $(\mathcal{T}_{2},\mathcal{D}_{2})$. This cascaded design ensures that reliable high-confidence detections are prioritized, while low-confidence detections are selectively used to recover missed targets without introducing excessive false positives.
IV Approach
To improve data association in multi-object tracking, we build upon the two-stage matching strategy introduced in ByteTrack (see Background III). In this scheme, detections are divided into high-confidence and low-confidence groups: the former provide reliable initial associations, while the latter are selectively used to recover missed targets. This design enhances recall in crowded or ambiguous scenes without substantially increasing false positives. Throughout the paper, we follow the variable notation established in the Background, and introduce additional variables when describing our own extensions.
While effective, this two-stage association alone is insufficient under severe occlusion and dense traffic, where matching ambiguities frequently lead to identity switches or drift. Moreover, the naive initialization and deletion policies used in the baseline and ByteTrack can further degrade performance, highlighting the need for more selective strategies. To address these shortcomings, our framework augments the baseline with three key components: (1) explicit occlusion detection metrics based on spatial overlap, allowing occluded targets to be identified and handled even without detections; (2) environment-aware constraints—such as road directionality, street layout, and pedestrian zones—that refine motion trajectories and prevent implausible movements; and (3) revised initialization and deletion policies that reduce spurious identities while improving recovery of true targets. Together, these additions significantly improve robustness in complex, multi-class urban scenes. The complete FastTrack algorithm is summarized in Figure 3 and algorithm below.
@@ IV Approach
Motion Prediction.   For each tracklet $t\in\mathcal{T}$, we estimate its future state $\hat{x}_{t}$ using a class-aware Kalman filter, $\mathrm{KalmanPredict}(t)$. The motion model parameters are selected based on the object class: vehicles such as cars or motorcycles are allowed higher velocities and acceleration bounds, whereas pedestrians are modeled with smoother and slower dynamics. This enables more realistic trajectory predictions, especially in cases of temporary occlusion or detection dropout.
@@
Direction and ROI Constraint.   To prevent tracklets from drifting into implausible directions, we retrieve a region $R$ around the predicted state $\hat{x}_{t}$ via a region lookup on the environment map $\mathcal{M}$. Each region encodes semantic layout (e.g., drivable roads, pedestrian paths) and is associated with a motion cone that specifies the allowed directionality. The cone’s opening angle $\theta$ is derived from the polygonal structure of $R$: given an entrance edge and an exit edge (defined by prior scene-flow knowledge), we construct the two crossing diagonals that connect opposite endpoints of the entrance and exit edges. Let the entrance edge endpoints are $E_{1},E_{2}$ and the exit edge endpoints are $O_{1},O_{2}$, with diagonals $\overrightarrow{E_{1}O_{2}}$ and $\overrightarrow{E_{2}O_{1}}$, then
$\theta\;=\;\arccos\!\left(\frac{(\,O_{2}-E_{1}\,)\cdot(\,O_{1}-E_{2}\,)}{\lVert O_{2}-E_{1}\rVert\,\lVert O_{1}-E_{2}\rVert}\right).$ (1)
For each active tracklet $t$, we also compute its instantaneous motion direction $\phi$. Let the track center positions at frames $k$ and kNk-N be $p_{k}=(x_{k},y_{k})$ and $p_{k-N}=(x_{k-N},y_{k-N})$. The displacement $\Delta p=p_{k}-p_{k-N}=(\Delta x,\Delta y)$ gives
$\phi\;=\;\mathrm{arctan}(2\Delta y,\Delta x).$ (2)
The function $\mathrm{ProjectToCone}(\hat{x}_{t},\mathrm{cone}(R))$ then projects the predicted position into the direction-constrained cone (i.e., enforces $\phi\in[-\theta/2,\theta/2]$ around the region’s dominant flow), and $\mathrm{ClampToROI}(\hat{x}_{t},R)$ keeps the state inside the region. We illustrate these concepts in Figure 4 for a one-way road; the same construction extends to two-way roads and crosswalks.
@@
Association.   In the first stage, high-confidence detections $\mathcal{D}_{\mathrm{high}}$ are matched with active tracklets $\mathcal{T}$ using IoU-based association, yielding matches $(\mathcal{T}_{1},\mathcal{D}_{1})$. Unmatched tracklets and low score detections, $\mathcal{T}_{\mathrm{remain}}$ and $\mathcal{D}_{\mathrm{low}}$, are then passed to the second stage, where $\mathcal{T}_{\mathrm{remain}}$ is matched with low-confidence detections $\mathcal{D}_{\mathrm{low}}$ to recover hard cases $\mathcal{T}_{2}$ and we seclude them from $\mathcal{T}_{\mathrm{remain}}$. Finally, the updated active set becomes $\mathcal{T}=\mathcal{T}_{1}\cup\mathcal{T}_{2}$.
@@
Occlusion Handling.   Unmatched tracklets $\mathcal{T}_{\mathrm{remain}}$ are checked for occlusion by measuring spatial overlap with active tracklets $\mathcal{T}$. If the center-proximity score $\mathrm{CP}(t,t^{\prime})$, calculated from the IoU, exceeds a threshold $\mathrm{CP}_{\min}$, the tracklet $t$ is considered occluded and added to $\mathcal{T}_{\mathrm{occ}}$. For each occluded tracklet, we mark it occluded and apply two corrective operators to stabilize its state during disappearance: (i) velocity damping and (ii) bounding-box enlargement, shown in Figure 5.
For velocity damping, we modify Kalman state of a tracklet represented as $x_{t}$ containing $(p_{x},p_{y})$ denote the position and $(v_{x},v_{y})$ the velocity. The operator $\mathrm{DampenVelocity}(t)$ reduces the velocity magnitude by a damping factor $\gamma_{\mathrm{velo}}\in(0,1)$:
$(v_{x},v_{y})\;\leftarrow\;\gamma_{\mathrm{velo}}\cdot(v_{x},v_{y}).$ (3)
In addition, to mitigate drift during occlusion, the position is softly reset toward its last non-occluded location:
$(p_{x},p_{y})\;\leftarrow\;(p_{x},p_{y})-\delta_{\mathrm{reset}}.$ (4)
This reset constrains unrealistic forward propagation, while subsequent Kalman updates continue from the corrected state.
For bounding-box enlargement, to improve re-identification after reappearance, the bounding box dimensions are slightly enlarged depending on the object class. If $(w,h)$ denote the width and height, then the operator $\mathrm{EnlargeBox}(t)$ applies
$(w,h)\;\leftarrow\;\beta_{\mathrm{enlarge}}\cdot(w,h).$ (5)
This adjustment increases tolerance for detector noise and partial visibility when the occluded object becomes visible again. Occluded tracklets $\mathcal{T}_{\mathrm{occ}}$ are then excluded from further association until reappearance, ensuring that identity drift is minimized. The effectiveness of velocity damping and bounding-box enlargement is illustrated in Figure 6.
Figure 5: Illustration of DampenVelocity and EnlargeBox over four frames (F=14F=1\!-\!4). Top: standard tracking, where fast motion and tight boxes cause drift and ID switches. Bottom: our method modifies $t_{2}$, where $t^{{}^{\prime}}_{2}\leftarrow t_{2}$ after DampenVelocity and $t^{{}^{\prime\prime}}_{2}\leftarrow t^{{}^{\prime}}_{2}$ after EnlargeBox, preserving consistent identities during occlusion.
@@ VI Experimental Setup
Implementation Details.    For detection, we adopt YOLOX due to its effective balance between speed and accuracy. The detector’s classification and localization heads are trained following the best practices established in prior work [5, 35]. At inference time, we apply a non-maximum suppression (NMS) threshold of 0.75. Tracklet association uses an IoU threshold of 0.5, and exponential moving average (EMA) smoothing is applied with a momentum coefficient $\alpha$ of 0.8. We set $\mathrm{CP}_{\min}$ to 0.7, the initialization overlap threshold $K_{init}$ to 0.8, and the occlusion tolerance window $T_{occ}$ to 30 frames. Detection confidence thresholds for classification are $\tau_{low}=0.2$ and $\tau_{high}=0.65$. The region of interest $\mathcal{M}$ is user-configurable, and if direction information is provided, the tracker adapts accordingly. We set $0.75\leq\gamma_{\mathrm{velo}}\leq 0.9$, $\delta_{\mathrm{reset}}\in\{3,4\}\mathrm{frames}$ and $1.1\leq\beta_{\mathrm{enlarge}}\leq 1.2$ depending on the object class type.
@@ VII-A Ablation Studies:
Post Processing: To evaluate the contribution of the post-processing stage, we analyze the impact of incorporating the GSI, which outperforms conventional interpolation techniques [37], and global linking (G-Link) on the final tracking performance across three datasets. Importantly, all results reported here are based on the full online system, which includes our proposed deletion/initialization policies and occlusion handling mechanisms. The post-processing modules are applied offline after generating the initial tracks. As shown in Table V, enabling either GSP or G-Link individually brings consistent improvements over the base output, with G-Link slightly outperforming GSP in terms of HOTA across all datasets.
@@ VII-B Benchmark Evaluation
TABLE IX: MOT20.
FastTracker Benchmark:    Table X presents the results on the FastTracker Benchmark, a challenging internal benchmark designed to assess robustness in dense tracking scenarios. FastTracker achieves the best MOTA (63.8), IDF1 (79.2) and HOTA (61.0), surpassing FeatureSORT and ByteTrack in association accuracy. It also achieves the lowest number of identity switches (251), demonstrating strong consistency across frames. While FastTracker leads in MOTA (63.8), it also shows a better balance between detection and identity preservation, outperforming StrongSORT and FeatureSORT by +1.3 and +3 HOTA, respectively, and significantly improving IDF1 by up to +1.8. These results highlight FastTracker’s strength in maintaining accurate identity tracking under crowded conditions. Figure 8 illustrates the relationship between occlusion levels in 4 FastTracker benchmark videos and identity switches. The percentages shown in the bars denote the relative reduction in ID switches of FastTracker w.r.t. the baseline ByteTrack, indicating the outperformance of FastTracker. Notably, higher reductions are observed in videos with more occluded objects, highlighting the robustness of our method under challenging conditions.
@@
Figure 8: Reduction in ID switches vs. baseline. Bars: occluded objects (left axis). Lines: ID switches per method (right axis).
VIII Discussion
FastTracker operates in real time across all evaluated scenarios, sustaining frame-rate processing on the incoming streams even under severe occlusions and dense object/frame conditions. This efficiency enables deployment on a wide range of edge devices without sacrificing accuracy. While FastTracker demonstrates strong performance across public and internal benchmarks, some limitations remain. Currently, the system relies on manually defined ROI regions and cone direction constraints, which must be specified using exactly four edges. This configuration may limit deployment in complex or dynamic scenes where such annotations are impractical or insufficient. As a promising direction for future work, recent advances in semantic segmentation and scene understanding can be leveraged to enable automatic detection of road boundaries, crosswalks, and other contextual cues, removing the need for manual setup. Moreover, extending the system to support arbitrary polygonal ROIs or more flexible directional constraints would allow broader applicability in real-world environments such as intersections, roundabouts, and multi-lane roads.
IX Conclusion